OpenStudy (anonymous):

help please 7/10 x +6 = 1/2 x

OpenStudy (anonymous):

are you solving for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

.... would the answer be -7.4?

OpenStudy (anonymous):

i checked and apparently the answer was 30 i dont really understand how they get that though

OpenStudy (anonymous):

..... wowwww.... one sec let me try that again :)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

.... I got 7.4.... are you sure the answer is 30?

OpenStudy (anonymous):

yes im sure

OpenStudy (anonymous):

well.... i don't know what to do -.-... i feel like i'm not helping you at all.....

OpenStudy (anonymous):

i have another one... \[\frac{ 1 }{ 6} x +8 = \frac{ 1 }{ 2} x\]

OpenStudy (anonymous):

if that format helps a ltitle.. XD

OpenStudy (anonymous):

little*

OpenStudy (anonymous):

haha ok like they say if at first you don't succeed try try again... unfortunately whenever you use that piece of advice you usually become even more confused xD :)

OpenStudy (anonymous):

totally!! XD :)

OpenStudy (anonymous):

ok so when i plugged in 30 into the first equation i got this as the answer: 25.2=15 so the answer doesn't match up with the equations..... i don't know why.... >.<

OpenStudy (anonymous):

dang... eh ill just skip those fraction questions yeah that sounds good XD

OpenStudy (anonymous):

hahaha I wish you all the luck with those questions... i'm going to think about the question even when I don't have to..... my brain hurts XD lol :0

OpenStudy (anonymous):

lol :o XD

OpenStudy (anonymous):

i was trying to do the smiley face sign hahaha :) i hit the wrong key.... lol

OpenStudy (anonymous):

lol smooth cx jk :)

OpenStudy (anonymous):

thanksssss.... jk... hahahaha

OpenStudy (anonymous):

hahhahahaha cx

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)